OpenStudy (anonymous):

find all real solutions of x^3−7x^2+12x=0

OpenStudy (anonymous):

x^3-7x^2+12x = 0 x(x^2-7x+12) = 0 x[(x-3)(x-4)] = 0 x = 0, 3, 4

jhonyy9 (jhonyy9):

x(x2-7x+12)=0 x=0 x=3 x=4

OpenStudy (anonymous):

can any of you solve this one, x^4-x^3-42x^2=0, so far i have x^2(x^2-x-42)=0